# Market-Bench: Evaluating Large Language Models on Introductory Quantitative Trading and Market Dynamics

### v1

@@ Abstract
We introduce Market-Bench, a benchmark that evaluates large language models (LLMs) on introductory quantitative trading tasks by asking them to construct executable backtesters from natural-language strategy descriptions and market assumptions. Each instance specifies one of three canonical strategies—scheduled trading on Microsoft (NASDAQ: MSFT), pairs trading on Coca-Cola (NASDAQ: KO) and Pepsi (NASDAQ: PEP), or delta hedging on MSFT—and models must produce code whose P&L, drawdown, and position paths match a verifiable reference implementation. We assess twelve state-of-the-art models using a multi-round pass@k metric that separates structural reliability (whether the backtest runs) from numerical accuracy (mean absolute error of the backtest metrics). While most models reliably execute the simplest strategy (average pass@3 of $0.80$), errors vary by orders of magnitude across models and tasks: Gemini 3 Pro and Claude 4.5 Sonnet combine strong reliability with low error on simpler strategies, GPT-5.1 Codex-Max achieves perfect pass@1 on the first two strategies and the lowest best-run error on the easiest task, and Qwen3 Max attains perfect pass@3 yet sometimes produces inaccurate P&L paths. These results show that current LLMs can scaffold basic trading infrastructure but still struggle to reason robustly about prices, inventory, and risk; we release Market-Bench and a public leaderboard at: https://marketbench.ai.
@@ 2 Related Work
There has been increased work to build domain-specific large language models for finance. BloombergGPT, for example, is an early proprietary model trained on a large mixture of general and financial data to support tasks such as sentiment analysis, news classification, and question answering within the Bloomberg ecosystem (wu2023bloomberggpt). On the other hand, FinGPT proposes an open-source pipeline for financial large language models that emphasizes automatic data pulling and lightweight fine-tuning so that models can be continually adapted to new market information (yang2023fingpt). Furthermore, the PIXIU framework introduces an evaluation suite that covers multiple financial NLP and prediction tasks, providing one of the first publicly available financial LLM + benchmarks (xie2023pixiu). More recently, FinBen and Open-FinLLMs continue in this direction by covering a broad range of financial tasks and multimodal benchmarks and models (text, tables, time series, and charts), respectively (xie2024finben; xie2024openfinllms). In addition, FinanceQA also demonstrates that models fail 60% of realistic tasks at hedge funds and other financial institutions mateega2025financeqabenchmarkevaluatingfinancial. Most of the current work in this area evaluates high-level language tasks such as classification, extraction, and textual analysis.
@@
There have been several benchmarks on evaluating large language models in financial domains. FinEval targets Chinese financial domain knowledge through thousands of multiple-choice questions which span academic finance along with industry practice. In addition, CFinBench builds a comprehensive Chinese financial benchmark which tests professional qualification exams and roles like tax consultants and securities analysts (nie2024cfinbench). FinEval-KR further separates knowledge versus reasoning ability and introduces separate metrics and datasets to study both at the same time (dou2025finevalkr). BizFinBench evaluates practical, business driven applications like calculation, reasoning, and information extraction (lu2025bizfinbench). While these benchmarks provide broad coverage of financial reasoning and understanding, they do not test how these models could be practically used in a quantitative trader’s day to day life.
@@
On the code side, Market-Bench has similar aspects to general purpose code generation benchmarks. HumanEval evaluates models trained on code by asking them to create function bodies that pass unit tests (chen2021codex). MBPP (Mostly Basic Programming Problems) measures creation on short, natural language problems with unit test evaluation as well (austin2021programsynthesis). DS-1000 targets data science code generation in realistic settings by using problems from StackOverflow spanning seven Python libraries. It highlights the inconsistency of model generated code for data science tasks (lai2022ds1000). Furthermore, SWE-Bench tested models on 2,294 different software engineering problems jimenez2024swebenchlanguagemodelsresolve. Current state-of-the-art models can solve $\sim$70% of the latest SWE-bench problems. These benchmarks stress code correctness but are not applied to financial market structure. Market-Bench instead focuses on domain-specific backtesting and market mechanics, evaluating models both on code reliability and on error of P&L, positions, and risk metrics when compared to a verifiable implementation.
@@ 3.1 Data Collection
All of the datasets were either generated synthetically through a random process or obtained from (databento_us_equities). We preprocessed the datasets by randomizing the volume available at each price level and only using the top 3 levels available. This was done to ensure that the models tracked the liquidity that trades remove from the book and whether they persisted that liquidity correctly. Furthermore, the options delta dataset for Strategy 3 was generated using a simple random walk.
@@ 4.1 Per-Attempt Metrics
where $d$ is the number of scalar metrics (e.g., total P&L, max drawdown, etc.) produced by the backtester for the final state of the simulation. For FAILED attempts, MAE is undefined and not considered.
@@ 4.2 Per-Round Aggregation
$\text{solved}_{s,m,r}$ $=\begin{cases}1&\text{if at least one attempt is {SUCCESS}},\\ 0&\text{otherwise},\end{cases}$ (3)
$\text{first\_attempt}_{s,m,r}$ $=\min\{k:\text{attempt $k$ is {SUCCESS}}\}.$ (4)
If a round is never solved, we set solved to 0 and leave best_mae and first_attempt undefined. Models are ranked first on their consistency and then on their error.
@@ 4.3 Per-Strategy Metrics
$\text{pass@3}_{s,m}$ $=\frac{1}{5}\sum_{r=1}^{5}\text{solved}_{s,m,r},$ (5)
$\text{pass@1}_{s,m}$ $=\frac{1}{5}\sum_{r=1}^{5}\mathbf{1}\{\text{first\_attempt}_{s,m,r}=1\},$ (6)
$\text{mean\_mae\_solved}_{s,m}$ $=\frac{\sum_{r:\text{solved}_{s,m,r}=1}\text{best\_mae}_{s,m,r}}{\sum_{r=1}^{5}\text{solved}_{s,m,r}},$ (7)
$\text{avg\_attempt}_{s,m}$ $=\frac{\sum_{r:\text{solved}_{s,m,r}=1}\text{first\_attempt}_{s,m,r}}{\sum_{r=1}^{5}\text{solved}_{s,m,r}}.$ (8)
Here, pass@3 captures the fraction of rounds for which the model eventually produces at least one successful attempt. pass@1 captures how often the very first attempt suffices. mean_mae_solved measures numerical fidelity on solved rounds, and avg_attempt measures how many attempts it typically takes to reach a successful backtest output.
@@ 4.4 Overall Metrics and Ranking
To obtain an overall view across strategies, we average each metric over $s\in\{1,2,3\}$:
$\text{overall\_pass@3}_{m}$ $=\frac{1}{3}\sum_{s=1}^{3}\text{pass@3}_{s,m},$ (9)
$\text{overall\_pass@1}_{m}$ $=\frac{1}{3}\sum_{s=1}^{3}\text{pass@1}_{s,m},$ (10)
and define overall mean MAE and average attempts as the mean of the per-strategy values over strategies where the model solves at least one round.
Within each strategy, we rank models by:
  1. 1.
    Higher pass@3,
  2. 2.
    Higher pass@1,
  3. 3.
    Lower mean_mae_solved,
  4. 4.
    Lower avg_attempt.
This prioritizes reliability (solving rounds) before accuracy, and accuracy before efficiency in number of attempts.
All scalar figures reported in the tables and text below (pass@k, MAE, and average attempts) are rounded to two decimal places.
@@ 5.1 Strategy-Level Results
Tables 13 show the per-strategy rankings for all models, sorted from best to worst. Each table lists pass@3, pass@1, mean MAE on solved rounds, best run MAE, and average attempt index to first success.
@@ 5.1.1 Strategy 1: Single-Stock Scheduled Execution
Model pass@3 pass@1 Mean MAE (solved) Best run MAE Avg. attempts
gemini-3-pro-preview 1.00 1.00 14.83 14.83 1.00
claude-sonnet-4.5 1.00 1.00 16.36 16.35 1.00
mistral-large-2512 1.00 1.00 361.87 23.01 1.00
gpt-5.1-codex-max 1.00 1.00 844.74 0.002 1.00
llama-4-maverick 1.00 1.00 4,137.62 170.06 1.00
deepseek-v3.2 1.00 0.80 133.63 7.26 1.40
qwen3-max 1.00 0.80 2,388.18 16.39 1.20
grok-4 0.80 0.00 59.35 7.22 2.25
claude-opus-4.5 0.60 0.40 14.10 7.18 1.33
llama-3.1-nemotron-ultra 0.60 0.40 8,799.23 111.63 1.67
nova-premier-v1 0.40 0.20 688.57 171.13 1.50
command-a 0.20 0.00 630.21 630.21 3.00
Table 1: Strategy 1: Scheduled execution on MSFT. Seven of twelve models solve all five rounds (pass@3 = 1.00).
Strategy 1 is the easiest task. No model fails all five rounds, and the average pass@3 across all models is $0.80$, with five of the models having perfect pass@1. However, the errors of the outputs vary considerably.
  • GPT-5.1 Codex-Max achieves the lowest best-run MAE (0.002) with perfect pass@3 and pass@1, though its mean MAE (844.74) is higher due to variance across rounds.
  • Gemini 3 Pro and Claude Sonnet combine reliability (pass@3 = 1.00, pass@1 = 1.00) with very small mean MAE (14.83 and 16.36, respectively).
  • DeepSeek V3.2 and Mistral-Large-2512 also solve all rounds but with higher MAE (133.63 and 361.87), while Llama-4 Maverick has substantially larger error (4,137.62) despite perfect pass@3 and pass@1.
  • Qwen3 Max, Amazon Nova Premier, Nvidia Nemotron, and Cohere Command-A frequently produce executable backtests but with errors in the hundreds to thousands.
@@ 5.1.2 Strategy 2: Pairs Mean-Reversion on COKE/PEP
Model pass@3 pass@1 Mean MAE (solved) Best run MAE Avg. attempts
gemini-3-pro-preview 1.00 1.00 52.22 52.22 1.00
gpt-5.1-codex-max 1.00 1.00 136.97 89.43 1.00
claude-sonnet-4.5 1.00 0.80 205.36 70.82 1.20
mistral-large-2512 1.00 0.80 267.21 135.24 1.40
qwen3-max 1.00 0.60 572,587,991.97 100.14 1.40
grok-4 0.80 0.40 573.78 119.25 1.75
deepseek-v3.2 0.60 0.60 132.10 125.87 1.00
llama-4-maverick 0.40 0.20 3,202.11 131.92 2.00
llama-3.1-nemotron-ultra 0.40 0.20 14,010.61 135.26 2.00
claude-opus-4.5 0.40 0.00 138.40 85.56 2.50
command-a 0.40 0.00 7,335.00 6,738.19 3.00
nova-premier-v1 0.00 0.00
@@
Strategy 2 is more structurally challenging. Amazon Nova Premier never produces a successful attempt (pass@3 = 0.00), and several other models solve only two of the five rounds. On average, models achieve pass@3 of 0.67 and pass@1 of 0.47. The errors of the model outputs also show a wide spread:
  • Gemini 3 Pro again stands out, with perfect pass@3 = 1.00, pass@1 = 1.00, and the lowest mean MAE of 52.22.
  • GPT-5.1 Codex-Max and DeepSeek V3.2 achieve low MAE (136.97 and 132.10, respectively), though DeepSeek solves only three of five rounds.
  • Claude Sonnet and Mistral-Large-2512 both solve all rounds with moderate MAE (205.36 and 267.21).
  • Qwen3 Max attains perfect pass@3 but has an MAE on the order of $5.73\times 10^{8}$, reflecting extreme divergence in logic from the verifiable backtester.
@@ 5.1.3 Strategy 3: Delta Hedging with MSFT
Model pass@3 pass@1 Mean MAE (solved) Best run MAE Avg. attempts
grok-4 1.00 1.00 1,482.33 1,013.33 1.00
claude-sonnet-4.5 1.00 1.00 16,157.31 805.16 1.00
qwen3-max 1.00 1.00 329,700.43 1,087.64 1.00
gemini-3-pro-preview 1.00 0.80 1,245.48 1,013.99 1.20
gpt-5.1-codex-max 1.00 0.80 8,326.53 1,370.72 1.20
deepseek-v3.2 1.00 0.60 12,075.80 1,127.14 1.40
mistral-large-2512 0.80 0.40 137,773.85 1,351.55 2.00
claude-opus-4.5 0.40 0.40 8,143.48 1,370.72 1.00
command-a 0.40 0.20 21,488.70 19,852.90 1.50
llama-4-maverick 0.20 0.00 20,418.58 20,418.58 2.00
nova-premier-v1 0.00 0.00
llama-3.1-nemotron-ultra 0.00 0.00
@@
Strategy 3 is the most complex and numerically the harshest task. Two models (Amazon Nova Premier and Nvidia Nemotron) never solve a round (pass@3 = 0.00), while six models achieve pass@3 = 1.00. Average pass@3 across models is 0.65, and average pass@1 is 0.52.
  • Gemini 3 Pro achieves the lowest mean MAE (1,245.48) with pass@3 = 1.00 and pass@1 = 0.80.
  • Grok 4 also performs well, with perfect pass@3 and pass@1 and a mean MAE of 1,482.33.
  • GPT-5.1 Codex-Max, Claude Sonnet, and DeepSeek V3.2 solve all rounds but with MAE in the range of 8,326.53–16,157.31.
@@
  • Llama-4 Maverick, Mistral-Large-2512, Cohere Command-A, and Claude Opus failed to reason or implement the prompt correctly and had large MAEs.

### v2

@@ Abstract
We introduce Market-Bench, a benchmark that evaluates large language models (LLMs) on introductory quantitative trading tasks by asking them to construct executable backtesters from natural-language strategy descriptions and market assumptions. Each instance specifies one of three canonical strategies—scheduled trading on Microsoft (NASDAQ: MSFT), pairs trading on Coca-Cola (NASDAQ: KO) and Pepsi (NASDAQ: PEP), or delta hedging on MSFT—and models must produce code whose P&L, drawdown, and position paths match a verifiable reference implementation. We assess thirteen state-of-the-art models using a multi-round evaluation that separates structural reliability (whether the backtest runs) from numerical accuracy (mean absolute error of the backtest metrics), assigning failed outputs a duplicated-metrics baseline MAE. While most models reliably execute the simplest strategy (average Executable Passes of $4.08$ out of $5$ rounds), errors vary by orders of magnitude across models and tasks: Gemini 3 Pro and Claude 4.5 Sonnet combine strong reliability with low error on simpler strategies, GPT-5.2 achieves strong overall performance with perfect executability, GPT-5.1 Codex-Max achieves the lowest best-run error on the easiest task, and Qwen3 Max attains perfect executability yet sometimes produces inaccurate P&L paths. These results show that current LLMs can scaffold basic trading infrastructure but still struggle to reason robustly about prices, inventory, and risk; we release Market-Bench and a public leaderboard at: https://marketbench.ai.
@@ 2 Related Work
There has been increased work to build domain-specific large language models for finance. BloombergGPT, for example, is an early proprietary model trained on a large mixture of general and financial data to support tasks such as sentiment analysis, news classification, and question answering within the Bloomberg ecosystem (Wu and others, 2023). On the other hand, FinGPT proposes an open-source pipeline for financial large language models that emphasizes automatic data pulling and lightweight fine-tuning so that models can be continually adapted to new market information (Yang et al., 2023). Furthermore, the PIXIU framework introduces an evaluation suite that covers multiple financial NLP and prediction tasks, providing one of the first publicly available financial LLM + benchmarks (Xie and others, 2023). More recently, FinBen and Open-FinLLMs continue in this direction by covering a broad range of financial tasks and multimodal benchmarks and models (text, tables, time series, and charts), respectively (Xie and others, 2024a; b). In addition, FinanceQA also demonstrates that models fail 60% of realistic tasks at hedge funds and other financial institutions Mateega et al. (2025). Most of the current work in this area evaluates high-level language tasks such as classification, extraction, and textual analysis.
@@
There have been several benchmarks on evaluating large language models in financial domains. FinEval targets Chinese financial domain knowledge through thousands of multiple-choice questions which span academic finance along with industry practice. In addition, CFinBench builds a comprehensive Chinese financial benchmark which tests professional qualification exams and roles like tax consultants and securities analysts (Nie and others, 2024). FinEval-KR further separates knowledge versus reasoning ability and introduces separate metrics and datasets to study both at the same time (Dou and others, 2025). BizFinBench evaluates practical, business driven applications like calculation, reasoning, and information extraction (Lu and others, 2025). While these benchmarks provide broad coverage of financial reasoning and understanding, they do not test how these models could be practically used in a quantitative trader’s day to day life.
@@
On the code side, Market-Bench has similar aspects to general purpose code generation benchmarks. HumanEval evaluates models trained on code by asking them to create function bodies that pass unit tests (Chen and others, 2021). MBPP (Mostly Basic Programming Problems) measures creation on short, natural language problems with unit test evaluation as well (Austin and others, 2021). DS-1000 targets data science code generation in realistic settings by using problems from StackOverflow spanning seven Python libraries. It highlights the inconsistency of model generated code for data science tasks (Lai and others, 2022). Furthermore, SWE-Bench tested models on 2,294 different software engineering problems Jimenez et al. (2024). Current state-of-the-art models can solve $\sim$70% of the latest SWE-bench problems. These benchmarks stress code correctness but are not applied to financial market structure. Market-Bench instead focuses on domain-specific backtesting and market mechanics, evaluating models both on code reliability and on error of P&L, positions, and risk metrics when compared to a verifiable implementation.
@@ 3.1 Data Collection
All of the datasets were either generated synthetically through a random process or obtained from (Databento Inc., 2025). We preprocessed the datasets by randomizing the volume available at each price level and only using the top 3 levels available. This was done to ensure that the models tracked the liquidity that trades remove from the book and whether they persisted that liquidity correctly. Furthermore, the options delta dataset for Strategy 3 was generated using a simple random walk.
@@ 4.1 Per-Attempt Metrics
where $d$ is the number of scalar metrics (e.g., total P&L, max drawdown, etc.) produced by the backtester for the final state of the simulation.
For attempts that do not produce comparable outputs (failed attempts), we assign a strategy specific baseline MAE by repeating the initial (first row) metrics for all timestamps and computing its MAE against the true metrics time series. This baseline corresponds to not building a backtest at all as it repeats what was initially given.
@@ 4.2 Per-Round Aggregation
and if a round is never solved, we set best_mae to the strategy-specific baseline MAE and treat the round as unsolved for executability counting.
@@ 4.3 Per-Strategy Metrics
$\text{Mean MAE}_{s,m}$ $=\frac{1}{5}\sum_{r=1}^{5}\text{best\_mae}_{s,m,r},$ (3)
$\text{Best Run MAE}_{s,m}$ $=\min_{r\in\{1,\dots,5\}}\text{best\_mae}_{s,m,r},$ (4)
$\text{Executable Passes}_{s,m}$ $=\sum_{r=1}^{5}\mathbf{1}\{\text{at least one attempt is {SUCCESS}}\}.$ (5)
Here, Mean MAE averages numerical error across all rounds, including baseline-filled failures; Best Run MAE captures the best numerical fidelity achieved in any round; and Executable Passes counts how many rounds produced an executable backtest.
@@ 4.4 Overall Metrics and Ranking
To obtain an overall view across strategies, we aggregate per model over all $15$ (strategy, round) combinations:
$\text{Mean MAE}_{m}$ $=\frac{1}{15}\sum_{s=1}^{3}\sum_{r=1}^{5}\text{best\_mae}_{s,m,r},$ (6)
$\text{Best Run MAE}_{m}$ $=\min_{s\in\{1,2,3\},\,r\in\{1,\dots,5\}}\text{best\_mae}_{s,m,r},$ (7)
$\text{Executable Passes}_{m}$ $=\sum_{s=1}^{3}\sum_{r=1}^{5}\mathbf{1}\{\text{at least one attempt is {SUCCESS}}\}.$ (8)
Within each strategy (and overall), we rank models by: Mean MAE. All scalar figures reported in the tables and text below (MAE and executable passes) are rounded to two decimal places.
@@ 5.1 Strategy-Level Results
Tables 13 show the per-strategy rankings for all models, sorted from best to worst. Each table lists mean MAE, best run MAE, and executable passes.
@@ 5.1.1 Strategy 1: Single-Stock Scheduled Execution
Model Mean MAE Best Run MAE Executable Passes
gemini-3-pro-preview 14.83 14.83 5
claude-sonnet-4.5 16.36 16.35 5
claude-opus-4.5 111.40 7.18 3
grok-4 121.85 7.22 4
deepseek-v3.2 139.19 7.26 5
command-a 184.90 153.10 1
nova-premier-v1 242.34 153.10 2
mistral-large-2512 361.87 23.01 5
gpt-5.1-codex-max 844.74 0.002 5
gpt-5.2 1,545.89 153.10 5
qwen3-max 2,015.66 16.39 5
llama-3.1-nemotron-ultra 2,511.13 111.63 3
llama-4-maverick 4,137.62 170.06 5
Table 1: Strategy 1: Scheduled execution on MSFT. Eight of thirteen models solve all five rounds (Executable Passes = 5).
Strategy 1 is the easiest task. No model fails all five rounds, and the average Executable Passes across all models is $4.08$ out of $5$. However, the errors of the outputs vary considerably.
  • GPT-5.1 Codex-Max achieves the lowest best-run MAE (0.002) with Executable Passes = 5, though its mean MAE (844.74) is higher due to variance across rounds.
  • Gemini 3 Pro and Claude Sonnet combine reliability (Executable Passes = 5) with very small mean MAE (14.83 and 16.36, respectively).
  • DeepSeek V3.2 and Mistral-Large-2512 also solve all rounds but with higher MAE (139.19 and 361.87), while Llama-4 Maverick has substantially larger error (4,137.62) despite Executable Passes = 5.
  • Qwen3 Max, GPT-5.2, Amazon Nova Premier, Nvidia Nemotron, and Cohere Command-A frequently produce executable backtests but with errors in the hundreds to thousands.
@@ 5.1.2 Strategy 2: Pairs Mean-Reversion on COKE/PEP
Model Mean MAE Best Run MAE Executable Passes
gemini-3-pro-preview 52.22 52.22 5
gpt-5.2 107.02 48.10 5
deepseek-v3.2 132.77 125.87 3
nova-premier-v1 133.10 133.10 0
claude-opus-4.5 133.86 85.56 2
gpt-5.1-codex-max 136.97 89.43 5
claude-sonnet-4.5 193.32 70.82 5
mistral-large-2512 228.90 133.10 5
grok-4 309.37 119.25 4
llama-4-maverick 605.26 131.92 2
command-a 1,093.36 133.10 2
llama-3.1-nemotron-ultra 2,268.11 133.10 2
qwen3-max 408,991,460.86 100.14 5
@@
Strategy 2 is more structurally challenging. Amazon Nova Premier never produces a successful attempt (Executable Passes = 0), and several other models solve only two of the five rounds. On average, models achieve Executable Passes of 3.46. The errors of the model outputs also show a wide spread:
  • Gemini 3 Pro again stands out, with Executable Passes = 5 and the lowest mean MAE of 52.22.
  • GPT-5.2 and GPT-5.1 Codex-Max achieve low MAE (107.02 and 136.97, respectively), and both have Executable Passes = 5.
  • Claude Sonnet and Mistral-Large-2512 both solve all rounds with moderate MAE (193.32 and 228.90).
  • Qwen3 Max attains Executable Passes = 5 but has an MAE on the order of $4.09\times 10^{8}$, reflecting extreme divergence in logic from the verifiable backtester.
@@ 5.1.3 Strategy 3: Delta Hedging with MSFT
Model Mean MAE Best Run MAE Executable Passes
gpt-5.2 1,369.59 1,365.42 5
grok-4 1,482.33 1,013.33 5
gemini-3-pro-preview 4,595.52 1,013.99 5
gpt-5.1-codex-max 10,496.40 1,370.72 5
deepseek-v3.2 14,724.36 1,127.14 5
claude-sonnet-4.5 16,157.31 805.16 5
claude-opus-4.5 18,945.34 1,370.72 2
llama-4-maverick 21,279.52 20,418.58 1
nova-premier-v1 21,345.75 21,345.75 0
llama-3.1-nemotron-ultra 21,345.75 21,345.75 0
command-a 21,369.58 19,852.90 2
mistral-large-2512 63,683.24 1,351.55 4
qwen3-max 329,700.43 1,087.64 5
@@
Strategy 3 is the most complex and numerically the harshest task. Two models (Amazon Nova Premier and Nvidia Nemotron) never solve a round (Executable Passes = 0), while seven models achieve Executable Passes = 5. Average Executable Passes across models is 3.38, and mean MAE spans several orders of magnitude.
  • GPT-5.2 achieves the lowest mean MAE (1,369.59) with Executable Passes = 5.
  • Grok 4 also performs well, with Executable Passes = 5 and a mean MAE of 1,482.33.
  • GPT-5.1 Codex-Max, Claude Sonnet, and DeepSeek V3.2 solve all rounds but with MAE in the range of 10,496.40–16,157.31.